OpenStudy (anonymous):

(5w to the power of 2 - 7w -8) - (7w to the power of 2 + 3w -6) Simplify The Difference. Pleaasse Help! =)

OpenStudy (anonymous):

-2w^2-10w-2

OpenStudy (anonymous):

^2 Means To The Power Of, Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks Pal, It Means A Lot, thank You =)

OpenStudy (anonymous):

you can simplify more

OpenStudy (anonymous):

-w^2-5w-1

OpenStudy (anonymous):

\[(5w^2-7w-8)-(7w^2+3x-6)\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Yeaahs

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

remove () using by distributing the "-" sign, that is take opposite of everything inside the ( ) \[5w^2-7w-8-7w^2-3w+6\] then combine like terms:\[5w^2-7w^2=-2w^2\] \[-7w-3w=-10w\] \[-8+6=-2\] answer: \[-2w^2-10w-2\]

OpenStudy (anonymous):

you cannot simplify this further.

OpenStudy (anonymous):

Yay! Thank You! <3

OpenStudy (anonymous):

if you divide by 2 you will get the sipmlest form